Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 352     Entry time: 22 Sep 2023, 06:51
Author: Benjamin van Kuiken 
Group: SCS 
Subject: Shift summary 

Achievemnets

  • Characterized the X-ray and laser danage on the STO2LNO2

Problems

  • Cound not find fine timeing from XAS measurements performed on the sample
    • MCP sees 266 nm PP laser
    • SiN mesaurement not successful so far. It is possible that there is an issue with the SN membranes on the sample holder

Plan for next shift

  • Find fine timing
  • conclude damad sutdy of sample.
ELOG V3.1.4-7c3fd00